OpenStudy (anonymous):

find the exact value of the expression for the given value of theta. do not use a calculator. cos(theta over 2) if theta=5 pi over 3

OpenStudy (anonymous):

\[\theta =5\pi/3 \implies \theta/2 = 5\pi/6 \]

OpenStudy (anonymous):

This is not a difficult angle to do.

OpenStudy (anonymous):

You could use: \[ \cos(\theta/2)= \pm\sqrt{\frac{1+\cos(\theta)}{2}} \]The \(\pm\) is determined by the quadrant.